OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 + 4x3 + 6x2 + 4x + 5; d(x) = x2 + 1 f(x) = (x2 + 1)( x2 + 4x + 5) + 12x - 15 f(x) = (x2 + 1)( x2 + 4x + 5) f(x) = (x2 + 1)( x2 - 4x + 5) f(x) = (x2 + 1)( x2 - 4x + 5) + 12x - 15

OpenStudy (anonymous):

f(x) = (x2 + 1)( x2 + 4x + 5)

OpenStudy (anonymous):

\[\frac{x^4+4 x^3+6 x^2+4 x+5}{x^2+1}=x^2+4 x+5 \text{ and no remainder} \]